Shift summary for Owl 11/8/99 (fwd)

ugur@cuhep.Colorado.EDU
Mon, 8 Nov 1999 06:33:57 -0700 (MST)

Owl Shift, Nov. 8 1999

On Shift: S. Ledovskoy, U. Erkarslan

Hours of beam available at all intensities: 7.0 hours
Hours of beam available at nominal intensity: 6.6 hours
Hours of E799 data (special runs not included): 6.0 hours
Hours of special runs (laser scans, muon runs...): 0.1 hours

Accounting of lost time:Tape changing, Tevatron problem,drt_fresh_start

========================================================================
Runs ended this shift:

14923 BEAM_799 16.5 million events
14924 MUON_RUN
14925 BEAM_799 O.45 million events
=======================================================================
Tapes taken to FCC this shift

none

========================================================================
Accesses during shift:
none
========================================================================
Overview of shift: E799 datataking, Muon run, drt_fresh_start

========================================================================

Description of Shift:

14923 01:14 online split request

04:07 Ktevana error

05:43 Tapes are full and L3 acceptance fraction warning.
Changing tapes. Ramping down NM4AN(pos->neg). changing
NMSR polarity

14924 06:07 start taking triggers in muon run

06:10 stop muon run. Ramping up NM4AN

14925 06:22 Start BEAN_799 run

06:40 DAQ dead due to DDD veto 1. stop run. try to
drt_fresh_start
Init aborts (tape drives problem)
Plan is to replace tapes, do fresh start and clean
tapedrives in ktev1 ... in progress
we took out tapes *12-14, *38which have muon run and
15 min of data

========================================================================
Follow up Notes on Previous Shift